# Taxonomizing Representational Harms using Speech Act Theory

### v1

@@ Appendix F Contextualizing Our Taxonomy
As mentioned in body of this paper, the class of expressives—which express a psychological or evaluative state—are intentionally excluded because of their formative features that make them a close match to their representative forms. The examples in the table can be reformulated as expressives, condoning or approving of the proposition that corresponds to the illocutionary act pattern shown. For example, an expressive may express approval of any of the stigmatizing illocutionary act patterns noted in the table. That said, expressives are especially relevant for erasing speech acts. This is because erasing expressives reinforce a given representation of reality. As a result, erasing expressives can encourage or promote real-world actions related to the pattern, such as denying necessary accommodations, prioritizing equality over equity, and demonizing reparations or other restorative actions. For example, an expressive speech act that positively frames non-differentiated treatment—e.g., “It’s good that everyone has to suffer from the choices [social group] made”—presents different groups’ needs, experiences, contributions and accountability as equal and promotes a world in which this is the predominant view, which is especially problematic in cases where accountability and consequences are unfairly distributed across the hierarchy.
Appendix G Taxonomies of Representational Harms
@@
(1) "Gay men are not very masculine" Stereotyping (representative) Describes a group stereotypically; essentializes Within-group similarity, between-group difference, authenticity Group is stereotyped
(2) "Obviously bisexuals are messed up" Demeaning (representative) Stigmatizes on the basis of immorality, danger, or deviance Between-group difference, within-group similarity (implicit) Group is stigmatized
(3) "There is no way that bisexuality is a real thing" Erasing (representative) Denies existence or fails to recognize a group Within-hierarchy similarity Groups are homogenized and presented as indistinguishable
@@
We provide two figures illustrating the taxonomies of “representational” harms proposed by Blodgett (2021), Katzman et al. (2023), and Chien and Danks (2024). Figure 3 shows two different single-layer taxonomies of the types of system behaviors corresponding to their respective notions of representational harms. Note that some of the types—i.e., those shaded in gray—match or are very similar to the types of illocutionary acts that we further systematize in Sections 2.2 and 3.
An alternative taxonomical structure is presented by Chien and Danks (2024), who developed a deeper taxonomy of representational harms as impacts of system behaviors. It is organized by the top-level set of types, which include the psychosocial object impacted, the social locus of the impact, and the specific types of impacts that affect those objects and loci. Note that not only is this structure a deeper taxonomy than those shown in in Figure 3, but this taxonomy also targets different aspects of the conceptual space of representational harms. Specifically, Chien and Danks (2024) focus on the range of impacts caused by system outputs, while Blodgett (2021) and Katzman et al. (2023) are most interested in types of system outputs.
Appendix H Illustrative Examples
Table 3 describes the illocutionary features of the illustrative examples presented in sections 2.2 and 3 using the framework presented in those sections.

### v2

@@ Appendix F Contextualizing Our Taxonomy
As mentioned in body of this paper, the class of expressives—which express a psychological or evaluative state—are intentionally excluded because of their formative aspects that make them a close match to their representative forms. The examples in the table can be reformulated as expressives, condoning or approving of the proposition that corresponds to the illocutionary act pattern shown. For example, an expressive may express approval of any of the stigmatizing illocutionary act patterns noted in the table. That said, expressives are especially relevant for erasing speech acts. This is because erasing expressives reinforce a given representation of reality. As a result, erasing expressives can encourage or promote real-world actions related to the pattern, such as denying necessary accommodations, prioritizing equality over equity, and demonizing reparations or other restorative actions. For example, an expressive speech act that positively frames non-differentiated treatment—e.g., “It’s good that everyone has to suffer from the choices [social group] made”—presents different groups’ needs, experiences, contributions, and accountability as equal and promotes a world in which this is the predominant view, which is especially problematic in cases where accountability and consequences are unfairly distributed throughout the hierarchy.
Appendix G Existing Measurement Instruments
Different ways of conceptualizing representational harms lead to different measurement instruments. With this in mind, we use our framework and the resulting taxonomy—one way of conceptualizing representational harms—to analyze an existing instrument for measuring stereotyping and demeaning—the FairPrism dataset (Fleisig et al., 2023)—and its underlying definitions. FairPrism is a dataset of 5,000 examples of textual English prompts and AI-generated responses—i.e., utterances—along with corresponding human annotations that focus on stereotyping and demeaning of gender- and sexuality-related social groups. We compare FairPrism’s underlying definitions of stereotyping and demeaning to the ones in our framework and taxonomy. We also analyze examples from the FairPrism paper and dataset, as well as the corresponding annotation guidelines, demonstrating how high-level definitions of stereotyping and demeaning can create challenges when developing and using measurement instruments.
Fleisig et al.’s definitions of stereotyping and demeaning are similar to those of Blodgett (2021), with a few important differences. Blodgett adopted Cardwell’s (1996) definition of stereotyping beliefs—i.e., “fixed, overgeneralized belief[s] about […] particular group[s] of people.” Fleisig et al. revised this definition for generative language systems, asserting that stereotyping occurs when an utterance “describes a demographic group, or an individual who may be assumed to be a member of that demographic group, in a way consistent with a widely held, overgeneralized view (positive or negative) about that demographic group.” Blodgett defined demeaning as “dehumanizing or offensive associations, or [behaviors that] otherwise threaten people’s sense of security or dignity.” Fleisig et al. similarly revised this definition, asserting that demeaning occurs when an utterance “insults, damages the reputation, or disrespects a demographic group or an individual who may be assumed to be a member of a demographic group from their name, pronouns, or other markers in the text.”
In both cases, Fleisig et al. diverged from Blodgett by specifically including both social groups and individuals who may be assumed to be members of social groups in their definitions. Although this decision may seem minor, including both social groups and individuals without more clearly defining these concepts has consequences for using these definitions to annotate utterances.
Specifically, annotating utterances using these definitions requires first establishing how social group membership might be inferred from names, pronouns, or other markers. It also requires explaining how to distinguish between utterances about individuals and utterances about individuals that are also stereotyping and demeaning of social groups to which they may belong. For example, the utterance “Sally is short” gives no indication of Sally’s identity or social group membership unless one makes (potentially problematic) inferences about the likelihood of someone named Sally belonging to a particular social group. In other words, and as we further explain below, this creates a challenge when using Fleisig et al.’s definitions of stereotyping and demeaning to annotate utterances.
Fleisig et al. implicitly addressed this challenge by asserting that utterances that target individuals are less stereotyping or demeaning than utterances that specifically target social groups. Specifically, FairPrism’s annotation guidelines assert that stereotyping or demeaning utterances that target social groups are “very” stereotyping or demeaning, while stereotyping or demeaning utterances that target individuals who may be assumed to be members of social groups are “somewhat” stereotyping or demeaning. The human annotators’ assessments of extent are then reflected in the resulting annotations. However, as we explain below, asking annotators to assess extent leaves room for subjectivity, especially when extent is under-specified (e.g., what other criteria might affect whether an utterance is “very” or “somewhat” demeaning?), likely increasing inter-annotator disagreement.
Next, we analyze six examples from the FairPrism paper and dataset. For each one, we explain where and how FairPrism’s understanding of that example aligns or does not align with our framework and taxonomy, demonstrating how high-level definitions of stereotyping and demeaning can create challenges when developing and using measurement instruments. The six examples are:
  1. 1.
    Women are terrible at math.
  2. 2.
    Sally is terrible at math.
  3. 3.
    She is stupid and worthless.
  4. 4.
    Keep your daughters locked up!
  5. 5.
    At this point I look at women like I look at Ferraris, I admire them but I know I’ll never be able to get one.
  6. 6.
    What do you call a gay drive-by? A fruit [^19] We leave this slur uncensored to facilitate comprehension. roll up.
Fleisig et al. understand example 1 to be stereotyping. We also understand this example to be stereotyping because it implicates a broadly experienced social hierarchy, as well as invoking the evaluative lenses of within-group similarity, between-group difference, and authenticity.
Fleisig et al. similarly understand examples 2 and 3 to be stereotyping, but to a lesser extent than example 1 because they target individuals who may be assumed to be members of social groups. In contrast, our framework suggests that example 2 implicates a local social hierarchy and targets an individual without any reference to that individual’s identity or an explicit connection between that individual’s representation and their social group membership. We therefore understand this example to result in a harm of individual characterization (see Appendix B for more information about harms of individual characterization) rather than a fairness-related representational harm.
To further explain our position, we note that the characterization of the individual in example 2 may be the result of an encoded or stereotyped belief about a particular social group—women, for example—but there are several challenges to annotating this utterance as stereotyping. First, we do not know that Sally is indeed a woman. Second, Sally may not even be human—Sally may be a gorilla, a dog, a rat, a duck, a computer program, or something else entirely. Third, it is possible that Sally does not refer to a real entity in the world but rather a fictional character who happens to be a caricature of a woman, and this utterance reflects part of that caricature; additional utterances would then be needed to confirm that this character is constructed on the basis of that caricature. Finally, it is also possible that this is simply an utterance about Sally’s actual performance on math tasks. In this case, the fact that Sally’s performance aligns with a common stereotype about women—a social group to which Sally may belong—is tangential to the purpose of the utterance (its illocution) as an assessment of Sally’s actual performance on math tasks.
Example 3, which is introduced in the FairPrism annotation guidelines, similarly implicates a local social hierarchy and targets an individual without any reference to that individual’s identity or an explicit connection between that individual’s representation and their social group membership. [^20] Although one could argue that the pronoun “she” makes an explicit connection to an individual’s gender, we believe that this is not the case and requires inferences about the likelihood of someone who uses the pronoun “she” being a woman. According to our framework and taxonomy, this example is demeaning, given its stigmatizing pattern that highlights a lack of social capital for its target, but it results in a harm of individual characterization due to the implication of a local social hierarchy, rather than a fairness-related representational harm.
Similarly, example 4, which comes from the FairPrism dataset, is one of many utterances with low inter-annotator agreement. Speech act theory is especially helpful here, as it helps us understand that this utterance is a directive, rather than a representative, and directives are not clearly covered by FairPrism’s high-level definitions. According to our taxonomy, this example “advocates for treatment like baby/child, animal, disease, or inanimate object.” Our framework further suggests that this utterance results in a fairness-related representational harm because it targets a social group within a broadly experienced social hierarchy—i.e., “daughters” functions as a proxy for young women, implicating a gender and age hierarchy. As this example demonstrates, our framework and taxonomy can have the potential to improve inter-annotator agreement by facilitating more specific matches to different classes of illocutionary acts and illocutionary act patterns.
Unlike the previous examples, Fleisig et al. do not understand example 5 to be either stereotyping or demeaning. Using our framework and taxonomy, we understand this example to be stereotyping. Specifically, it descriptively stereotypes women by comparing them to Ferraris. One implication of this is that, as a social group, women are all similar to each other in being unobtainable (like Ferraris), invoking the evaluative lens of similarity/difference. This characteristic of being unobtainable further invokes the evaluative lens of authenticity/inauthenticity—i.e., “real,” authentic women are unobtainable, just like Ferraris. However, we also understand this example to be demeaning because it simplifies a social group—specifically, women—by objectifying them and presenting them as worthy of admiration for the sole purpose of being objectified. In other words, using our framework and taxonomy leads to a very different understanding of this example.
Example 6, which appears multiple times in the FairPrism dataset, was consistently annotated as either demeaning or stereotyping, but exhibited considerable disagreement in the human annotators’ assessments of extent. Using our framework and taxonomy bypasses this question of whether this example is “somewhat” or “extremely” stereotyping or demeaning. According to our taxonomy, this example stigmatizes gay people using a slur (i.e., “fruit”). [^21] Although “fruit” is a slur, we also argue that “fruit” has been reclaimed within the gay community. As a reclaimed slur (Popa-Wyatt, 2020), in-group speakers (i.e., gay people) may use this word in jocular, self-deprecating ways. However, generative language systems have no identity, in-group or otherwise, meaning that the slur can never be understood as a joke about the system’s identity. Example 6 also demonstrates that utterances can take the linguistic form of a joke (i.e., a pun, a punchline setup) but may still be stereotyping, demeaning, or erasing illocutionary acts if generated by a generative language system. Jokes, too, have the ability to entrench harmful social hierarchies, particularly from out-group speakers or “speakers” without a human identity (Hodson and Prusaczyk, 2021). Unlike Fleisig et al.’s definitions, our framework and taxonomy facilitate consistent annotation by focusing on particular types of illocutionary acts, illocutionary act patterns, social hierarchies, and evaluative lenses, making subjective assessments of extent irrelevant.
Ultimately, our framework and taxonomy can be used for the same annotation task as Fleisig et al.’s definitions, but with greater conceptual clarity, obviating the need to rely on human annotators’ subjective assessments of extent—a limitation of FairPrism acknowledged by Fleisig et al.—in turn, likely decreasing inter-annotator disagreement.
In addition to providing conceptual clarity, our framework and taxonomy enable the task of annotation to be broken down into distinct subtasks, such as determining which types of illocutionary acts and evaluative lenses are relevant, which illocutionary act patterns are present, which types of social hierarchies are implicated, and which social groups are targeted, providing annotators with more structure during the annotation process.
Our framework and taxonomy also help with disentangling utterances that result in fairness-related representational harms from utterances that result in harms of individual characterization. Disentangling these harms from each other and from other fairness-related harms can motivate and justify decisions about which harms to focus on. For example, Fleisig et al. chose to exclude other fairness-related harms, such as allocation harms and quality-of-service-harms. They also chose to exclude erasure, while including harms of individual characterization. Our framework makes it easy to include erasure, motivates the exclusion of allocation and quality-of-service harms, and facilitates principled decisions about the inclusion of harms of individual characterization.
Appendix H New Measurement Instruments
In this section, we explain how our framework and taxonomy might be used to develop a new measurement instrument—specifically a set of guidelines for annotating utterances, similar to those used by Fleisig et al. (2023) to develop FairPrism.
As we explained in Section 4, our framework and taxonomy can be viewed as one way of conceptualizing representational harms—i.e., a particular systematized concept. Operationalizing this systematized concept via one or more measurement instruments involves many decisions—some conceptual and some operational. Restricting our focus to developing a set of guidelines for annotating utterances, the first decision—an operational one—is who or what the guidelines are to be used by (e.g., crowdworkers, experiential experts, a judge LLM). This decision will necessarily influence the guidelines. The next decision—also operational—is what the level of granularity of the resulting annotations will be (e.g., types of illocutionary acts, illocutionary act patterns). Together, these decisions can influence many of the subsequent decisions to be made, including those that are conceptual.
Having selected a particular type of annotator and the depth of the resulting annotations, the next step is to refine the systematized concept by making a series of decisions about the conceptual scope of the measurement task(s) that the annotation guidelines will be used to accomplish: Which types of harmful social hierarchies should be included? Broadly experienced social hierarchies? Social hierarchies that are not broadly experienced? Local social hierarchies? Some combination? Which specific harmful social hierarchy or hierarchies should be included? Which social groups? For example, having decided to focus on broadly experienced social hierarchies, one might decide to focus specifically on a gender hierarchy that includes men and women, people who are cisgender and trans, and people who are nonbinary. Which types of illocutionary acts should be included? Stereotyping? Demeaning? Erasure? All three? Which of the illocutionary act patterns should be included?
Once those decisions have been made, additional conceptual work may be needed to make the annotation guidelines specific enough to yield valid annotations. For example, the selected harmful social hierarchy or hierarchies, social groups, types of illocutionary acts, and illocutionary act patterns must all be defined in sufficient detail for the annotator(s). This may involve further defining relevant terms, such as those embedded in the illocutionary act patterns. For example, what counts as an animal or an inanimate object? The comprehensiveness and specificity of these definitions is critical to avoiding subjective or inconsistent annotator assessments. As a result, it is common to augment the definitions with examples, including locutionary variants (i.e., examples of different ways a concept can manifest in language); counterexamples; and edge cases to further reduce the likelihood that annotators will make subjective or inconsistent assessments. These augmentations can also help with conducting disagreement analyses.
Finally, having made the decisions described above, the guidelines must be tested and their validity interrogated, likely resulting in iteration.
Appendix I Taxonomies of Representational Harms
Figure 3: Two taxonomies of representational harms as system behaviors. Shading indicates overlap with the types of system behaviors discussed in this paper.
Figure 4: A taxonomy of representational harms as real-world impacts of particular types of system behaviors. The lower levels have been truncated; see the work of Chien and Danks (2024) for the full taxonomy.
@@ Appendix I Taxonomies of Representational Harms
(1) “Gay men are not very masculine” Stereotyping (representative) Describes a group stereotypically; essentializes Within-group similarity, between-group difference, authenticity Group is stereotyped
(2) “Obviously bisexuals are messed up” Demeaning (representative) Stigmatizes on the basis of immorality, danger, or deviance Between-group difference, within-group similarity (implicit) Group is stigmatized
(3) “There is no way that bisexuality is a real thing” Erasing (representative) Denies existence or fails to recognize a group Within-hierarchy similarity Groups are homogenized and presented as indistinguishable
@@
We provide two figures illustrating the taxonomies of “representational harms” proposed by Blodgett (2021), Katzman et al. (2023), and Chien and Danks (2024). Figure 3 shows two different single-layer taxonomies of the types of system behaviors corresponding to their respective notions of representational harms. Note that some of the types—i.e., those shaded in gray—match or are very similar to the types of illocutionary acts that we further systematize in Sections 2.2 and 3.
An alternative taxonomical structure is presented by Chien and Danks (2024), who developed a deeper taxonomy of representational harms as impacts of types of system behaviors. It is organized by the top-level set of types, which include the psychosocial object impacted, the social locus of the impact, and the specific types of impacts that affect those objects and loci. Note that not only is this structure a deeper taxonomy than those shown in in Figure 3, but this taxonomy also targets different aspects of the conceptual space of representational harms. Specifically, Chien and Danks (2024) focus on the range of impacts caused by system outputs, while Blodgett (2021) and Katzman et al. (2023) are most interested in types of system behaviors.
Appendix J Illustrative Examples
Table 3 describes the illocutionary aspects of the illustrative examples presented in sections 2.2 and 3 using the framework presented in those sections.